OpenStudy (anonymous):

Let f(x) = 7x - 13. Find f-1(x).

OpenStudy (anonymous):

@Hero

hero (hero):

Let y = f(x), then y = 7x - 13 swap x and y: x = 7y - 13 Now isolate y

OpenStudy (anonymous):

so x + 13 =7y

hero (hero):

Okay, but you still have one step left in order to isolate y.

hero (hero):

In order for y to be isolated, it has to be alone on one side.

OpenStudy (anonymous):



OpenStudy (anonymous):

like that?

hero (hero):

Okay, yes

OpenStudy (anonymous):

then I don't get what's after that

hero (hero):

What happens if you ever have 7/7?

OpenStudy (anonymous):

well then it's canceling each other out

hero (hero):

Exactly. The last step is to replace y with \(f^{-1}(x)\)

OpenStudy (anonymous):

so it's